# VLG-CBM: Training Concept Bottleneck Models with Vision-Language Guidance

### v1

@@ Abstract
Concept Bottleneck Models (CBMs) provide interpretable prediction by introducing an intermediate Concept Bottleneck Layer (CBL), which encodes human-understandable concepts to explain models’ decision. Recent works proposed to utilize Large Language Models (LLMs) and pre-trained Vision-Language Models (VLMs) to automate the training of CBMs, making it more scalable and automated. However, existing approaches still fall short in two aspects: First, the concepts predicted by CBL often mismatch the input image, raising doubts about the faithfulness of interpretation. Second, it has been shown that concept values encode unintended information: even a set of random concepts could achieve comparable test accuracy to state-of-the-art CBMs. To address these critical limitations, in this work, we propose a novel framework called Vision-Language-Guided Concept Bottleneck Model (VLG-CBM) to enable faithful interpretability with the benefits of boosted performance. Our method leverages off-the-shelf open-domain grounded object detectors to provide visually grounded concept annotation, which largely enhances the faithfulness of concept prediction while further improving the model performance. In addition, we propose a new metric called Number of Effective Concepts (NEC) to control the information leakage and provide better interpretability. Extensive evaluations across five standard benchmarks show that our method, VLG-CBM, outperforms existing methods by at least 4.27% and up to 51.09% on accuracy at NEC=5, and by at least 0.45% and up to 29.78% on average accuracy across different NECs, while preserves both faithfulness and interpretability of the learned concepts as demonstrated in extensive experiments [^1] Our code will be available at https://github.com/Trustworthy-ML-Lab/VLG-CBM.
Figure 1: We compare the decision explanation of VLG-CBM with existing methods by listing top-5 contributions for their decisions: (1) VLG-CBM provides concise and accurate concept attribution for the decision; (2) LF-CBM[13] frequently uses negative concepts for explanation, which is less informative; (3) LM4CV[18] attributes the decision to concepts that do not match the images, a reason for this is that LM4CV uses a limited number of concepts, which hurts CBM’s ability to explain diverse images; (4) Both LF-CBM and LM4CV have a significant portion of contribution from non-top concepts, making decisions less transparent. Full figure is in Appendix Fig. D.1.
@@ 1 Introduction
As deep neural networks become popular in real-world applications, it is crucial to understand the decision of these black-box models. One approach to provide interpretable decisions is the Concept Bottleneck Model(CBM) [4], which introduced an intermediate concept layer to encode human-understandable concepts and make final predictions based on these concept predictions. Unfortunately, this approach requires dense concept annotations from human experts, making it expensive and less applicable in practice as labels may not always be available.
@@
    Inaccurate concept prediction: Their concept predictions often contain factual errors i.e. the predicted concepts do not match the image. Moreover, as concepts are generated by LLMs, there are some non-visual concepts, for example "loud music" or "location" used in LF-CBM[13], which further hurt the faithfulness of concept prediction.
@@
    Information leakage: Recently, [11, 10] observed the information leakage in CBMs – the concept prediction encodes unintended information for downstream tasks, even if the concepts are irrelevant to the task.
In this paper, we propose a new framework called Vision-Language-Guided Concept Bottleneck Model (VLG-CBM) to address these two challenges. Our contributions are summarized below:
@@
    We propose to use the open-domain grounded object detection model to generate localized, visually recognizable concept annotations. This approach automatically filters the non-visual concepts. Furthermore, the location information is utilized to augment the data.
@@
    We provide the first theoretical explanations for information leakage in CBMs. Building on our theory, we propose a new metric called the Number of Effective Concepts (NEC), which facilitates fair comparison between different CBMs. We also show that using NEC can help to effectively control information leakage and enhance interpretability in our VLG-CBM.
@@
    We conduct a series of experiments and demonstrate that our VLG-CBM outperforms existing methods across 5 standard benchmarks by at least 4.27% and up to 51.09% on accuracy at NEC=5, and by at least 0.45% and up to 29.78% on average accuracy across different NECs. Our learned CBM achieves a high sparsity of 0.2% in the final layer even on large datasets including Places365, preserving interpretability even with a large number of concepts. Additionally, we qualitatively demonstrate that our method provides more accurate concept attributions compared to existing methods.
@@
    LF-CBM $\triangle$ $\checkmark$ $\checkmark$ $\triangle$ $\times$ $\triangle$
    Labo $\times$ $\checkmark$ $\times$ $\triangle$ $\times$ $\triangle$
    LM4CV $\checkmark$ $\times$ $\times$ $\triangle$ $\triangle$ $\triangle$
@@ 2 Related work
Concept Bottleneck Model (CBM). The Concept Bottleneck Model [4] introduces an intermediate concept bottleneck layer (CBL), where each neuron represents a human-understandable concept. CBL is followed by a linear prediction layer, which maps concepts to classes, enabling interpretable final decisions. Formally, let feature representation generated by a frozen backbone represented by $z=\phi(x)$, CBL concept prediction as $g(z)=W_{c}z$, and the final prediction layer as $h(\cdot)=W_{F}g(z)+b_{F}$. The final class prediction of the CBM is given by $\hat{y}=h(g(z))=h\circ g\circ\phi(x)$.
@@
Under this setting, the key in training a CBM is obtaining an annotated {(image, concept)} paired dataset for training concept bottleneck layer $g$. [4] used human-specified labels to train the CBL in a supervised way. However, obtaining labels with human annotators is tedious and costly. Recently, [13], [18], and [19] proposed to utilize Large Language Models (LLM) to generate a set of concepts $S$, then train CBL by aligning image and concepts with the guidance of vision language models (e.g. CLIP). [13] proposed LF-CBM to train CBM by directly learning a mapping from the embedding space of backbone to concept values in the CLIP space using cosine cubed loss function, and then mapping concepts to classes using sparse linear layer. [18] proposed LM4CV, a task-guided concept searching method that learns text embeddings in the CLIP space, and then maps the learned embeddings to concepts obtained from LLM using nearest neighbor. [19] proposed LaBo, using submodular optimization to reduce the concept set, followed by using CLIP backbone for obtaining concept values. However, as shown in Sections 5.1 and 5.3 , these methods suffer from multiple issues: (1) The concept prediction is often incorrect and does not capture the visual attributes required for downstream class prediction (2) VLMs like CLIP suffer from modality gap between image and text embeddings [8] resulting in encoding unintended information, and even random concepts can achieve high accuracy [18]. We hypothesize that grounding the concepts on the training dataset using an open-domain object detection model and then using the obtained concepts for learning CBL leads to a more faithful representation of fine-grained concepts and avoids the modality gap issues introduced by VLMs. Table 1 demonstrates the superiority of VLG-CBM over existing methods [13, 18, 19] on properties including controlling information leakage, flexibility to use any backbone and accurate concept prediction.
@@
Open Domain Language Grounded Object Detection. Recent works including GLIP [7], GLIPv2 [21], and GroundingDINO [9] provide a method to detect objects in images by conditioning on natural language queries. In this work, we propose to use these methods for generating automatically-labeled concept dataset for training CBMs. This removes the need for human labelers, which is costly, tedious, and does not scale to large datasets.
@@
Figure 2: Pipeline depicting our automated Language-Vision Guided approach for generating concept dataset for training Concept Bottleneck Models.
@@ 3.1 Automated generation of auxiliary dataset
  • Vision supervision from Open-domain Object Detectors to ground candidate concepts to spatial information: We propose using recent advancements in Grounded Object Detection [9, 6, 21] for obtaining spatial localization of candidate concepts in the dataset. Particularly, we use Grounding-DINO[9] Swin-B model for obtaining bounding boxes for candidate concepts. For each image $x_{i}$ with class label $c$ and candidate concepts $S_{c}$, we prompt Grounding DINO model with $S_{c}$ and obtain $K_{i}$ bounding boxes:
@@
After collecting bounding boxes for every image, we filter out the concepts that do not appear in any bounding box, and get our final concept set:
@@ 3.2 Training Concept Bottleneck Layer
After constructing the concept-labeled dataset $D^{\prime}$, we now define our approach to train the concept bottleneck layer for predicting the fine-grained concepts in the input image in a multi-label classification setting. Let $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be a backbone that generates $d$-dimensional embeddings $z=\phi(x)$ for input image $x$. Note that $\phi(x)$ can be a pre-trained backbone or trained from scratch. Define $g$ to be the Concept Bottleneck Layer (CBL) which maps embeddings to concept logits. We train a sequential CBM [4, 11] to predict concepts $\tilde{S}=g(\phi(x))$ in an image using Binary Cross Entropy (BCE) loss for multi-label prediction. Additionally, to improve the diversity of the concept-labeled dataset $D^{\prime}$, we augment the training dataset by cropping images to a randomly selected bounding box and modifying the target one-hot vector to predict the concept corresponding to the bounding box. Our optimization objective in terms of BCE loss can be written as:
@@ 3.3 Mapping Concept to Classes
In this section, we define our approach to training a sparse linear layer to obtain class labels from concepts in an interpretability-preservable manner. Let $h:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ be the sparse linear layer with weight matrix $W_{F}$ and bias $b_{F}$, which maps concept logits to class logits. We train the sparse layer using the original dataset $D$ by first obtaining concept logits from the trained CBL(frozen), normalizing the concept logits, and then using them to predict class logits. Our optimization objective in terms of Cross Entropy (CE) loss can be written as:
@@ 4 Unifying CBM evaluation with Number of Effective Concepts (NEC)
Besides training, another important challenge for CBM is: how to evaluate the semantic information learned in the CBL? Conventionally, the classification accuracy for final class labels is an important metric for evaluating CBMs, with the intuition that good classification accuracy indicates that we learn useful semantic information in the CBL. However, there may exist potential problems with this metric, as it has been shown that information leakage exists in jointly or sequentially trained CBM [11, 10]. That is to say, the CBL could contain unintended information that could be used for downstream classification, even if the concept is irrelevant to the task. In fact, recently [18] showed that, when increasing the number of concepts, a randomly selected concept set could approach the accuracy of the concept set chosen with sophistication, supporting the existence of information leakage.
To better understand this phenomenon, in section 4.1, we conduct a first theoretical analysis to investigate random CBL and its capability. To the best of our knowledge, this is the first formal analysis of random CBL. Then, inspired by our theoretical result, we propose a new evaluation metric for CBM, named NEC in section 4.2.
@@ 4.1 Theoretical analysis of the Random CBL
We start by defining the notations. Denote $k$ the number of concepts in CBL. We assume that the CBL $g$ consists of a single linear layer: $g(z)=W_{c}z$, where $W_{c}\in\mathbb{R}^{k\times d}$ and $z\in\mathbb{R}^{d}$, and the final layer $h$ is also linear: $h\circ g(z)=W_{F}g(z)+b_{F}$, where $W_{F}\in\mathbb{R}^{C\times k}$, $b_{F}\in\mathbb{R}^{C}$. This is the common setting for CBMs. The following theorem suggests a surprising conclusion: a linear classifier upon random (i.e. untrained) CBL could accurately approximate any linear classifier trained directly on the representation, as the number of concepts in the CBL goes up. {thm}[] Suppose $\Sigma\in\mathbb{R}^{d\times d}$ the variance matrix of the representation $z$ which is positive definite, $\lambda_{max}$ is the largest eigenvalue of $\Sigma$. The weight matrix $W_{c}\in\mathbb{R}^{k\times d}$ is sampled i.i.d from a standard Gaussian distribution. For any linear classifier which is built directly on the representation $z$, i.e. $f:\mathbb{R}^{d}\rightarrow\mathbb{R},f(z)=w^{\top}z+b$, could be approximated by another linear classifier $\tilde{f}$ on concept logits $g(z)=W_{c}z$, i.e. $f(z)\approx\tilde{f}(z)=\tilde{w}^{\top}g(z)+\tilde{b}$, with the expected square error upper-bounded by
@@
Here $E(k)=\mathbb{E}_{W_{c}}\left[min_{(\tilde{w},\tilde{b})}\mathbb{E}_{z}\;\left[|f(z)-\tilde{f}(z)|^{2}\right]\right]$ denotes the average square error, $w\in\mathbb{R}^{d}$, $\tilde{w}\in\mathbb{R}^{k}$.
@@
In Eq. 8, we consider a 1-D regression problem where we use a linear combination of concept bottleneck neurons to approximate any linear function. The multi-class classification result could be derived by applying Eq. 8 to each class logit (see Corollary A.1). From Eq. 8, we could see that the expected error goes down linearly when concept number $k$ increases, and achieves 0 when $k\geq d$, where $d$ is the dimension of backbone representation $z$. This suggests that, even with a random CBL (i.e. $W_{c}$ is simply drawn from a standard Gaussian distribution without training on any concepts), the classifier could still approximate the original classifier well and achieve good accuracy, when concept number $k$ is large enough. We defer the formal proof of Eq. 8 to Appendix A.

### v2

@@ Abstract
Concept Bottleneck Models (CBMs) provide interpretable prediction by introducing an intermediate Concept Bottleneck Layer (CBL), which encodes human-understandable concepts to explain models’ decision. Recent works proposed to utilize Large Language Models (LLMs) and pre-trained Vision-Language Models (VLMs) to automate the training of CBMs, making it more scalable and automated. However, existing approaches still fall short in two aspects: First, the concepts predicted by CBL often mismatch the input image, raising doubts about the faithfulness of interpretation. Second, it has been shown that concept values encode unintended information: even a set of random concepts could achieve comparable test accuracy to state-of-the-art CBMs. To address these critical limitations, in this work, we propose a novel framework called Vision-Language-Guided Concept Bottleneck Model (VLG-CBM) to enable faithful interpretability with the benefits of boosted performance. Our method leverages off-the-shelf open-domain grounded object detectors to provide visually grounded concept annotation, which largely enhances the faithfulness of concept prediction while further improving the model performance. In addition, we propose a new metric called Number of Effective Concepts (NEC) to control the information leakage and provide better interpretability. Extensive evaluations across five standard benchmarks show that our method, VLG-CBM, outperforms existing methods by at least 4.27% and up to 51.09% on accuracy at NEC=5, and by at least 0.45% and up to 29.78% on average accuracy across different NECs, while preserving both faithfulness and interpretability of the learned concepts as demonstrated in extensive experiments [^1] Our code is available at https://github.com/Trustworthy-ML-Lab/VLG-CBM.
Figure 1: We compare the decision explanation of VLG-CBM with existing methods by listing top-5 contributions for their decisions. Our observations include: (1) VLG-CBM provides concise and accurate concept attribution for the decision; (2) LF-CBM [15] frequently uses negative concepts for explanation, which is less informative; (3) LM4CV[25] attributes the decision to concepts that do not match the images, a reason for this is that LM4CV uses a limited number of concepts, which hurts CBM’s ability to explain diverse images; (4) Both LF-CBM and LM4CV have a significant portion of contribution from non-top concepts, making decisions less transparent. Full figure is in Appendix Fig. D.1.
@@ 1 Introduction
As deep neural networks become popular in real-world applications, it is crucial to understand the decision of these black-box models. One approach to provide interpretable decisions is the Concept Bottleneck Model (CBM) [6], which introduced an intermediate concept layer to encode human-understandable concepts. The model makes final predictions based on these concepts. Unfortunately, one major limitation of this approach is that it requires concept annotations from human experts, making it expensive and less applicable in practice as concept labels may not always be available.
@@
    Challenge #1: Inaccurate concept prediction. The concept predictions in these CBMs often contain factual errors i.e. the predicted concepts do not match the image. Moreover, as concepts are generated by LLMs, there are some non-visual concepts, for example "loud music" or "location" used in LF-CBM [15], which further hurt the faithfulness of concept prediction.
@@
    Challenge #2: Information leakage. Recently, [13, 12] observed the information leakage in CBMs through empirical experiments – they found that the concept prediction encodes unintended information for downstream tasks, even if the concepts are irrelevant to the task.
In this paper, we propose a new framework called Vision-Language-Guided Concept Bottleneck Model (VLG-CBM) to address these two major challenges. Our contributions are summarized below:
@@
    To address Challenge #1, we propose to use the open-domain grounded object detection model to generate localized, visually recognizable concept annotations in Section 3. This approach automatically filters the non-visual concepts. Furthermore, the location information is utilized to augment the data. As far as we know, our VLG-CBM is the first end-to-end pipeline to build CBM with vision guidance from open-vocabulary object detectors.
@@
    To address Challenge #2, we provide the first rigorous theoretical analysis which proves that CBMs have serious issues on information leakage in Section 4.1, whereas previous study on information leakage [12, 25] only provides empirical explanations. Building on our theory, we further propose a new metric called the Number of Effective Concepts (NEC) in Section 4.2, which facilitates fair comparison between different CBMs. We also show that using NEC can help to effectively control information leakage and enhance interpretability in our VLG-CBM.
@@
    We conduct a series of experiments in Section 5 and demonstrate that our VLG-CBM outperforms existing methods across 5 standard benchmarks by at least 4.27% and up to 51.09% on accuracy at NEC=5, and by at least 0.45% and up to 29.78% on average accuracy across different NECs. Our learned CBM achieves a high sparsity of 0.2% in the final layer even on large datasets including Places365, preserving interpretability even with a large number of concepts. Additionally, we qualitatively demonstrate that our method provides more accurate concept attributions compared to existing methods.
@@
    LF-CBM[15] $\triangle$ $\checkmark$ $\checkmark$ $\triangle$ $\times$ $\triangle$
    LaBo[27] $\times$ $\checkmark$ $\times$ $\triangle$ $\times$ $\triangle$
    LM4CV[25] $\checkmark$ $\times$ $\times$ $\triangle$ $\triangle$ $\triangle$
@@ 2 Related work
Concept Bottleneck Model (CBM). The seminal paper [1] first proposed self-explaining models by leveraging the idea of autoencoder to learn interpretable basis concepts in an unsupervised manner without pre-specified concepts. Later, [6] proposed to learn interpretable concepts with human specifications (labels) in the concept bottleneck layer (CBL), and coin the term Concept Bottleneck Models (CBM). CBL is followed by a linear prediction layer, which maps concepts to classes, enabling interpretable final decisions. Formally, let feature representation generated by a frozen backbone represented by $z=\phi(x)$, CBL concept prediction as $g(z)=W_{c}z$, and the final prediction layer as $h(\cdot)=W_{F}g(z)+b_{F}$. The final class prediction of the CBM is given by $\hat{y}=h(g(z))=h\circ g\circ\phi(x)$.
@@
Under this setting, the key in training a CBM is obtaining an annotated {(image, concept)} paired dataset for training concept bottleneck layer $g$. In [6], the authors used human-specified labels to train the CBL in a supervised way. However, obtaining labels with human annotators could be very tedious and costly. Recently, [15], [25], and [27] proposed to utilize Large Language Models (LLM) to generate a set of concepts $S$, then train CBL by aligning image and concepts with the guidance of vision language models (e.g. CLIP). For example, Oikarinen et al. [15] proposed LF-CBM to train CBM by directly learning a mapping from the embedding space of backbone to concept values in the CLIP space using cosine cubed loss function with the neuron interpretability tool[14], and then mapping concepts to classes using sparse linear layer. [25] proposed LM4CV, a task-guided concept searching method that learns text embeddings in the CLIP space, and then maps the learned embeddings to concepts obtained from LLM using nearest neighbor. Yang et al. [27] proposed LaBo, using submodular optimization to reduce the concept set, followed by using CLIP backbone for obtaining concept values. However, as we show in Sections 5.1 and 5.3 , these methods suffer from multiple issues: (i) The concept prediction is often incorrect and does not capture the visual attributes required for downstream class prediction (e.g. see Fig. 1 b )(ii) VLMs like CLIP suffer from modality gap between image and text embeddings [9] resulting in encoding unintended information, and even random concepts can achieve high accuracy [25]. To address these issues, we explicitly ground the concepts on the training dataset using an open-domain object detection model and then using the obtained concepts for learning CBL – this can ensure a more faithful representation of fine-grained concepts and avoids the modality gap issues introduced by VLMs. Table 1 demonstrates the superiority of VLG-CBM over existing methods [15, 25, 27] on properties including controlling information leakage, flexibility to use any backbone, and accurate concept prediction.
There are some recent works aim at addressing the challenges of CBMs. Similar to us, Pham et al. [16] uses an open-vocabulary object detection model to provide an explainable decision. However, their model is directly adapted from an OWL-ViT model, while our VLG-CBM uses an open-vocabulary object detection model to train a CBL over any base model, providing more flexibility. Additionally, their model requires pretraining to get best performance, while our VLG-CBM could be applied post-hoc to any pretrained model. Kim et al. [4] proposed to filter non-visual concepts by adding a vision activation term to the concept selection step, whereas VLG-CBM uses an open-vocabulary object detectors in multiple stage of CBM pipeline: for filtering non-visual concepts and the guiding training of concept bottleneck layer. Sun et al. [20] aims at eliminating the information leakage, and the authors evaluate the information leakage by measuring the performance drop speed after removing top-contributing concepts. This metric can be controlled by our proposed NEC metric, because the performance reach minimum after removing all contributing concepts. Roth et al. [17] demonstrate that random words and characters achieve comparable CLIP zero-shot performance on visual classification tasks. However, their work does not address information leakage problem and is a very different setting from our work. To date, most of the CBMs focused on vision domains, including this work. There are some recent work applying CBM approach to different domains and different tasks, e.g. interpretable language models for text classifications [21, 22, 11] and for continual learning [26]. We refer the interested readers to their papers for more details.
@@
Open Domain Language Grounded Object Detection. Recent works, including GLIP [8], GLIPv2 [29], and GroundingDINO [10] detect objects in images in an open-vocabulary manner conditioned on natural language queries. In this work we propose to utilize open-vocabulary object detectors for automatically generating grounded concept dataset for training CBMs. This removes the need for human labelers, which is costly, tedious, and does not scale to large datasets. Further, the detected objects provide necessary vision-guidance for CBMs training as demonstrated in our experiments.
@@
Figure 2: VLG-CBM  pipeline: We design automated Vision+Language Guided approach to train Concept Bottleneck Models.
@@ 3.1 Automated generation of auxiliary dataset
  • Vision supervision from Open-domain Object Detectors to ground candidate concepts to spatial information: We propose using Grounding-DINO[10] Swin-B, current state-of-the-art grounded object detector, for obtaining bounding boxes of candidate concepts in the dataset. For each image $x_{i}$ with class label $c$ and candidate concepts $S_{c}$, we prompt Grounding DINO model with $S_{c}$ and obtain $K_{i}$ bounding boxes:
@@
After collecting bounding boxes for every image, we filter out the concepts that do not appear in any bounding box, and get our final concept set $\tilde{S}$:
@@ 3.2 Training Concept Bottleneck Layer
After constructing the concept-labeled dataset $D^{\prime}$, we now define our approach to train the concept bottleneck layer for predicting the fine-grained concepts in the input image in a multi-label classification setting. Let $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be a backbone that generates $d$-dimensional embeddings $z=\phi(x)$ for input image $x$. Note that $\phi(x)$ can be a pre-trained backbone or trained from scratch. Define $g$ to be the Concept Bottleneck Layer (CBL) which maps embeddings to concept logits. We train a sequential CBM [6, 13] $g(\phi(x))$ to predict concepts in an image using Binary Cross Entropy (BCE) loss for multi-label prediction. Additionally, to improve the diversity of the concept-labeled dataset $D^{\prime}$, we augment the training dataset by cropping images to a randomly selected bounding box and modifying the target one-hot vector to predict the concept corresponding to the bounding box. Our optimization objective in terms of BCE loss can be written as:
@@ 3.3 Mapping Concept to Classes
In this section, we define our approach to training a sparse linear layer to obtain class labels from concepts in an interpretability-preservable manner. Let $h:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ be the sparse linear layer with weight matrix $W_{F}$ and bias $b_{F}$, which maps concept logits to class logits. We train the sparse layer using the original dataset $D$ by first obtaining concept logits from the trained CBL(frozen), normalizing the concept logits with the mean and variance on training set, and then using them to predict class logits. Our optimization objective in terms of Cross Entropy (CE) loss can be written as:
@@ 4 Unifying CBM evaluation with Number of Effective Concepts (NEC)
Besides training, another important challenge for CBM is: how to evaluate the semantic information learned in the CBL? Conventionally, the classification accuracy for final class labels is an important metric for evaluating CBMs, with the intuition that a good classification accuracy indicates that useful semantic information is learned in the CBL. However, purely using accuracy as the evaluation metric could be problematic, as it has been shown that information leakage exists in jointly or sequentially trained CBM [13, 12]. That is to say, the CBL could contain unintended information that could be used for downstream classification hence achieving high classification accuracy, even if the concept is irrelevant to the task. In fact, recently [25] showed that, when increasing the number of concepts, a randomly selected concept set could even approach the accuracy of the concept set chosen with sophistication, supporting the existence of information leakage.
To better understand this phenomenon, in section 4.1, we conduct a first theoretical analysis to investigate random CBL and its capability. To the best of our knowledge, this is the first formal analysis of random CBL. Next, inspired by our theoretical result, we propose a new evaluation metric for CBM, named NEC in section 4.2. NEC provides a way to control information leakage and enhance the interpretability of model decisions.
@@ 4.1 Theoretical analysis of the Random CBL
We start by defining the notations. Denote $k$ the number of concepts in CBL. We assume that the CBL $g$ consists of a single linear layer: $g(z)=W_{c}z$, where $W_{c}\in\mathbb{R}^{k\times d}$ and $z\in\mathbb{R}^{d}$, and the final layer $h$ is also linear: $h\circ g(z)=W_{F}g(z)+b_{F}$, where $W_{F}\in\mathbb{R}^{C\times k}$, $b_{F}\in\mathbb{R}^{C}$. This is the common setting for CBMs. The following theorem suggests a surprising conclusion: a linear classifier upon random (i.e. untrained) CBL could accurately approximate any linear classifier trained directly on the representation, as the number of concepts in the CBL goes up. {thm}[] Suppose $\Sigma\in\mathbb{R}^{d\times d}$ is the variance matrix of the representation $z$ which is positive definite, $\lambda_{max}$ is the largest eigenvalue of $\Sigma$, and the weight matrix $W_{c}\in\mathbb{R}^{k\times d}$ is sampled i.i.d from a standard Gaussian distribution. For any linear classifier $f$ which is built directly on the representation $z$, i.e. $f:\mathbb{R}^{d}\rightarrow\mathbb{R},f(z)=w^{\top}z+b$, it could be approximated by another linear classifier $\tilde{f}$ on concept logits $g(z)=W_{c}z$, i.e. $f(z)\approx\tilde{f}(z)=\tilde{w}^{\top}g(z)+\tilde{b}$, with the expected square error $E(k)$ upper-bounded by
@@
Here $E(k)=\mathbb{E}_{W_{c}}\left[\min_{(\tilde{w},\tilde{b})}\mathbb{E}_{z}\;\left[|f(z)-\tilde{f}(z)|^{2}\right]\right]$ denotes the average square error, $w\in\mathbb{R}^{d}$, $\tilde{w}\in\mathbb{R}^{k}$, $k$ is the number of concepts in CBL.
@@
In Eq. 8, we consider a 1-D regression problem where we use a linear combination of concept bottleneck neurons to approximate any linear function. The multi-class classification result could be derived by applying Eq. 8 to each class logit (see Corollary A.1). From Eq. 8, we could see that the expected error goes down linearly when concept number $k$ increases, and achieves 0 when $k\geq d$, where $d$ is the dimension of backbone representation $z$. This suggests that, even with a random CBL (i.e. $W_{c}$ is simply drawn from a standard Gaussian distribution without any training), the classifier could still approximate the original classifier well and achieve good accuracy, when concept number $k$ is large enough. We defer the formal proof of Eq. 8 to Appendix A.